1. The local Toyota dealer must decide how many tires of a particular type to order for replacement orders. This type of tire costs $25 each. The carrying charge is 10% per year and the ordering cost is $20 per order. Assume that the annual demand of the tires is 40000 units, the EOQ is:

A. 200 units

B. 219 units

C. 400 units

D. 800 units
Answer #1

Ordering cost (S) = $20 per order

Annual quantity demanded (D) = 40,000 units

Holding cost (H) = 25 x 10%

= $2.5

Economic order quantity = \sqrt{} (2SD)/H

= \sqrt{} (2 X 20 X 40,000)/2.5

= \sqrt{} 640,000/2.5

= 800 units

Kindly comment if you need further assistance. Thanks‼!